# Learning to Collaborate: A Capability Vectors-based Architecture for Adaptive Human-AI Decision Making

### v1

@@ 6.3.2 Random sampled decision makers
For changeable decision makers, the use of capability vector is essential as it is necessary to identify each decision maker’s capability. We assume that there are $M$ candidate decision makers, while each time only a subset of them participate in collaborative decision making. Consider the multiple classification task with 10 categories as for CIFAR10. We further assume that each human expert can have expertise in one to ten categories, which means there will be $10242^{10}=$ 1024 possible expert capabilities. Based on this assumption, we build expert vocabularies with different sizes $M$ by sampling the expertise categories (with replacement) as a binary vector across all possible expert capabilities (i.e. from 1024 combinations). For each training or evaluation instance, we sample different 2 experts from the expert vocabulary for both training and testing. We consider three settings: (a) Use learnable capability matrix as is shown in Figure 2; (b) Use fixed capability matrix; (c) Do not use capability vectors, just determine the two combination weights by the input image representations. We fixed the system random seed to 42 and compare different settings with the shared expert vocabulary. We apply an Adam optimizer with a learning rate of 1e-5 and the batch size equals to 64. The results are shown in Table 3.
@@
We notice that given the same model size, as the vocabulary size of decision makers becomes larger, the advantage of using learnable capability vectors becomes larger. Notice that the weight of capability matrix only takes less than 1% extra parameters of the weight generation model. This finding can be applied in building collaborative annotation systems with large number of candidate annotators. Moreover, the learning behavior of the capability vectors is investigated in Section 7.2.
@@ 6.4.1 Results on CIFAR10
In the first experiment, we follow (Mozannar and Sontag, 2020) to simulate multiple synthetic experts of varying competence: if the image belongs to the first $k$ classes with $k=1,2,...,10$, the expert predicts the correct label of the image, otherwise the expert predicts uniformly over all classes. We use a tunable resnet-18 pre-trained on CIFAR-10 as the feature extractor, and a 28-layer Wide-Resnet same with that in (Mozannar and Sontag, 2020) is applied as the AI classifier to be collaborative with human expert. To ensure that the AI classifier is the same for our method and the baseline method, we reproduce the baseline method in (Mozannar and Sontag, 2020) with 200 epochs of training, and then use the weights of the AI classifier to initialize the classifier of our proposed model. For training the whole architecture, we apply an Adam optimizer with a learning rate of 1e-5 and a batch size of 64, with 200 epochs of training. We take different levels of expert’s competence $k$, and compare the performance of our method with a set of baseline methods including Global (the proposed global weighted output merging methods in Section 5), Consistent (Mozannar and Sontag, 2020), Confidence (Raghu et al., 2019), OracleReject and MPZ18 (Madras et al., 2018). The results are given in Table  4. We can learn that our proposed architecture provide a consistent improvement from the baseline methods. In addition, there is no extra hyper-parameter involved in the loss function to be optimized. We also notice that if we extend the training epochs to 500 for the baseline method in (Mozannar and Sontag, 2020), gain of the system accuracy is not significant.
@@
Table 4: Comparison of system accuracy between our method and baselines with varying expert competence on CIFAR10. Among the baseline methods, for Consistent method, we reproduce the experiments by ourself. For other methods, we directly apply the reported results in (Mozannar and Sontag, 2020).
To further investigate the effect of human expert with relatively larger non-expertise capability level, we fix $k=5$ and change the non-expertise capability level by using different probability $p$. Using the same optimizer and hyper-parameter setting, the results are given in Table 5. We observe that our method achieve higher accuracy compared with the baseline models. Again, the proposed global weighted merging baseline achieves competitive performances compared with instance aware methods, but still outperformed by the proposed instance aware weighted merging architecture. In addition, we observe that as the non-expertise capability level increase, the baseline L2D method in (Mozannar and Sontag, 2020) has an accuracy drop when $p$ lies between 0.5 and 0.9. Meanwhile, the our proposed method, the system accuracy generally increases with the increase of non-expertise capability level. One possible explanation is that as the non-expertise capability level get close to the model accuracy, the baseline method may have instability in learning or be sensitive to hyper-parameter settings, which can be better resolved by merging the output of the model with human expert’s decision.
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
Ours 97.66 97.79 97.94 98.09 98.35 98.32 98.6 98.78 98.75 99.14 100
Global 96.46 96.71 96.97 97.25 97.54 97.85 98.18 98.53 98.91 99.36 100
Consistent 96.31 96.58 96.9 97.15 97.13 97.29 95.23 93.88 92.94 95.6 100
Table 5: Comparison of system accuracy between our method and baselines with non-expertise capability level on CIFAR10 with $k=5$.
@@
We further conduct an experiment on CIFAR 100, in which there are 100 classes and 20 super-classes of images. Instead of predicting the labels of fine-grained classes, we predict the labels of super-classes. We train a Wide-ResNet-28 on this classification task and get an accuracy of 73.33%, and utilize the pretrained model to initialize the AI classifier in both our proposed architecture and the baseline learning to defer method. We then compare our method with two alternative approaches:(a) Global, weighted output merging approach proposed in Section 5; (b) Consistent learning to defer estimator (Mozannar and Sontag, 2020), where each time the decision is made by only one expert or AI model. We tune the hyper-parameter of learning rates for both settings with a set of $\{1e^{-3},3e^{-4},1e^{-4},3e^{-5},1e^{-5},3e^{-6},1e^{-6}\}$. The results of varying expert competence is given in Table 6.
@@
It is shown that with the same initialized AI classifier and the experts’ competences, our method outperforms the baseline L2D method significantly in the cases of effective human-AI collaboration. Global weighted merging method achieve very competitive results as $k$ gets larger. Similary as in Section 6.4.1, we further investigate the effective of varying non-expertise capability levels with expert competence level $k=5$. The results are shown in Table 7. We observe that our method has a greatest advantage over the baseline around $p=0.7$, and for other non-expertise capability levels, it is possible that the advantage is smaller than the advantage at $p=0.0$.
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Ours 85.19 85.62 86.43 86.67 87.88 88.97 90.01 91.6 93.47 96.03
Global 77.27 78.95 80.75 82.67 84.70 86.83 89.06 91.42 93.94 96.66
Consistent 81.43 82.14 83.2 84.19 85.82 86.55 86.71 87.27 91.06 95.43
Table 7: Experiment on CIFAR100-superclass with different non-expertise capability levels.
@@ 6.4.3 Results on HateSpeech
We also conducted experiments on Hate speech dataset, in which we apply the proposed architecture and baseline methods to determine if an instance is hatespeech, offensive language or neither. We follow (Mozannar and Sontag, 2020) to create the synthetic experts for this task, which is: (a) If the tweet is in AAE then with probability $p$ the expert predict the correct label and otherwise predict uniformly at random. (b) If the tweet is not in AAE, the expert predict the correct label with probability $q$ and otherwise predict uniformly at random. Also, we consider three different expert probabilities $p$ and $q$, corresponding to a fair expert ($p=q=0.9$), a biased expert towards AAE ($p=0.75,q=0.9$), and a biased expert towards non AAE tweets ($p=0.9,q=0.75$). We use a pre-trained transformer encoder on Hatespeech to initialize our model as well as the classifier in the baseline model. We use a Adam optimizer with a learning rate of 1e-6 and a batch size of 32. The results are shown in Table 8. The collaborative method largely outperforms the model-only and human-only methods. Moreover, compared with the method of consistent estimator given in (Mozannar and Sontag, 2020), our method achieves significant improvements for non-AAE biased and fair setting.
@@
To investigate the effectiveness of the proposed architecture for the scenario with multiple human experts, we assume that we have multiple human experts with different expertise sets with equal size, for which each expert can have only 1 or 2 expertise categories. We consider the collaborative decision making by one model and these human experts. Also, we consider the case of solely using these experts for collaborative decision making without collaboration with AI model. The experiments are done on CIFAR10. We adopt a pretrained ResNet-18 model as the classifier as well as the image representation extractor. A 6 layer transformer encoder with 8 heads and a embedding dimension of 512 is applied. An Adam optimizer with a learning rate of 1e-5, and a batch size of 64 are applied. The result of the case that each expert has only 1 expertise category is shown in Table 9, where $s$ is the size of experts group. We observe that when the number of experts is small, the performance of both AI+experts system and experts-only system increase linearly. However, when the number of experts becomes large, the increasing trends diminishes slowly. This indicates that we need to apply heavy hyper-parameter tuning or use a more powerful weight generation model for handling the collaboration with a large number of experts.
0 1 2 3 4 5 6 7 8 9 10
AI+experts 93.52 95.82 96.76 97.21 97.2 97.3 97.29 97.34 97.27 97.28 97.38
Experts 0 19 25.78 31.21 35.68 39.31 41.96 45.78 48.97 52.77 56.12
Table 9: Comparison of system accuracy with varying number of experts.
In the second experiment of multiple expert collaborative decision making. we assume that each expert has an expertise set with size 2, and the 5 experts have expertise sets $\{1,2\}$, $\{3,4\}$, $\{5,6\}$, $\{7,8\}$, $\{9,10\}$, respectively. The result of this case that is shown in Table 10. Here We observe that similar trend is shown by the table. The propose method provide a significant improvement of the system accuracy as number of expert increases, but there is a diminishing marginal utility effect when the expert number is large.
0 1 2 3 4 5
AI+Experts 89.47 92.18 93.71 94.89 94.75 94.86
Experts 0 28 40.23 46.5 53.22 60.13
Table 10: Collaborative decision making with multiple experts. Each expert has 2 expertise categories.
@@ 7 Analysis
7.1 Effect of non-expertise capability levels.
We visualize the accuracy curves with different non-expertise correct probability for both the proposed method and the baseline consistent estimator in (Mozannar and Sontag, 2020), which is given by Figure 4. We can clearly observe that in both cases the baseline method has a breakdown of increasing trends when the non-expertise capability level is close to the accuracy of AI model. Especially on CIFAR10 dataset, the accuracy drop to about 93%-95% when the non-expertise capability level is between 0.6 and 0.9. In contrast, our proposed architecture show monotonic increase trends in both cases.
@@ 8.1 Comparing with related works
The proposed architecture adopts a strategy of merging the output vectors from both AI models and human experts, rather than merely deciding whether to defer the task to a human. In contrast to the work by (Kerrigan et al., 2021), which also focuses on output vector combination, our method incorporates a representation learning mechanism instead of relying on a naive Bayesian estimator based on a confusion matrix. This allows our approach to seamlessly accommodate multiple experts in an end-to-end manner.
While (Kerrigan et al., 2021) utilizes real human labels from the CIFAR-10H and ImageNet-16H datasets as the outputs from human experts, these datasets are annotated by a diverse group of human experts. Consequently, their estimation may not effectively adapt to the varying capabilities of different experts. Furthermore, their methodology rests on the assumption of conditional independence, which is often not satisfied in practice.
Our architecture also facilitates human-AI collaborative decision-making involving multiple AI models and multiple human experts. This is achieved through the learning of capability vectors in a uniform dimensional space. Related studies on human-AI collaborative decision-making with multiple experts typically employ different datasets or base models in their experiments. For instance, (Mao et al., 2024a) utilize ResNet-4 as the predictive AI model, while employing ResNet-10, ResNet-16, and ResNet-28 with increasing capacities as the human experts, focusing on the CIFAR-10 and SVHN image classification datasets. Additionally, Hemmer et al. (2022) investigate three distinct tasks, including Hate Speech detection, CIFAR-100 classification, and the NIH Dataset.
@@ 9 Conclusion
In this study, we present a human-AI collaborative decision-making architecture based on capability vectors, applicable to both generative and discriminative decision tasks. By utilizing capability vectors, we can uniformly model the decision intelligence of human experts and AI models across the same set of dimensions. Additionally, we developed a learning-free global weighted merging method for collaborative decision-making involving one human expert and one AI model specifically for classification tasks, which can be applied as a baseline for check the effectiveness of instance-aware weighted merging methods. For image classification tasks and sentiment analysis task, Our experiments, conducted on image classification and sentiment analysis tasks, demonstrate that the proposed method outperforms existing approaches across various human-AI collaborative scenarios. We show that for a widely used expert model in classification tasks, our method is more robust to the change of non-expertise capability level. In addition, the proposed method can be applied for multiple expert collaborative decision with or without AI decision makers.

### v2

@@ 5.3.2 Random sampled decision makers
For changeable decision makers, the use of capability vector is essential as it is necessary to identify each decision maker’s capability. We assume that there are $M$ candidate decision makers, while each time only a subset of them participate in collaborative decision making. Consider the multiple classification task with 10 categories as for CIFAR10. We further assume that each human expert can have expertise in one to ten categories, which means there will be $2^{10}=1024$ possible expert capabilities. Based on this assumption, we build expert vocabularies with different sizes $M$ by sampling the expertise categories (with replacement) as a binary vector across all possible expert capabilities (i.e. from 1024 combinations). For each training or evaluation instance, we sample different 2 experts from the expert vocabulary for both training and testing. We consider three settings: (a) Use learnable capability matrix as is shown in Figure 2; (b) Use fixed capability matrix; (c) Do not use capability vectors, just determine the two combination weights by the input image representations. We fixed the system random seed to 42 and compare different settings with the shared expert vocabulary. We apply an Adam optimizer with a learning rate of 1e-5 and the batch size equals to 64. The results are shown in Table 4.
@@
We notice that given the same model size, as the vocabulary size of decision makers becomes larger, the advantage of using learnable capability vectors becomes larger. Notice that the weight of capability matrix only takes less than 1% extra parameters of the weight generation model. This finding can be applied in building collaborative annotation systems with large number of candidate annotators. Moreover, the learning behavior of the capability vectors is investigated in Section 6.
@@ 5.4.1 Results on CIFAR10
In the first experiment, we follow (Mozannar and Sontag 2020) to simulate multiple synthetic experts of varying competence: if the image belongs to the first $k$ classes with k=1,2,,10k=1,2,...,10, the expert predicts the correct label of the image, otherwise the expert predicts uniformly over all classes. We use a tunable resnet-18 pre-trained on CIFAR-10 as the feature extractor, and a 28-layer Wide-Resnet same with that in (Mozannar and Sontag 2020) is applied as the AI classifier to be collaborative with human expert. To ensure that the AI classifier is the same for our method and the baseline method, we reproduce the baseline method in (Mozannar and Sontag 2020) with 200 epochs of training, and then use the weights of the AI classifier to initialize the classifier of our proposed model. For training the whole architecture, we apply an Adam optimizer with a learning rate of 1e-5 and a batch size of 64, with 200 epochs of training. We take different levels of expert’s competence $k$, and compare the performance of our method with a set of baseline methods including Global (the proposed global weighted output merging methods in Section 4), Consistent (Mozannar and Sontag 2020), Confidence (Raghu et al. 2019), OracleReject and MPZ18 (Madras et al. 2018). The results are given in Table  5. We can learn that our proposed architecture provide a consistent improvement from the baseline methods. In addition, there is no extra hyper-parameter involved in the loss function to be optimized. We also notice that if we extend the training epochs to 500 for the baseline method in (Mozannar and Sontag 2020), gain of the system accuracy is not significant.
Table 5: Comparison of system accuracy between our method and baselines with varying expert competence on CIFAR10. Among the baseline methods, for Consistent method, we reproduce the experiments by ourself. For other methods, we directly apply the reported results in (Mozannar and Sontag 2020).
@@
To further investigate the effect of human expert with relatively larger non-expertise capability level, we fix $k=5$ and change the non-expertise capability level by using different probability $p$. Using the same optimizer and hyper-parameter setting, the results are given in Figure 4 (a). We observe that our method achieve higher accuracy compared with the baseline models. Again, the proposed global weighted merging baseline achieves competitive performances compared with instance aware methods, but still outperformed by the proposed instance aware weighted merging architecture. In addition, we observe that as the non-expertise capability level increase, the baseline L2D method in (Mozannar and Sontag 2020) has an accuracy drop when $p$ lies between 0.5 and 0.9. Meanwhile, the our proposed method, the system accuracy generally increases with the increase of non-expertise capability level. One possible explanation is that as the non-expertise capability level get close to the model accuracy, the baseline method may have instability in learning or be sensitive to hyper-parameter settings, which can be better resolved by merging the output of the model with human expert’s decision.
@@
We further conduct an experiment on CIFAR 100, in which there are 100 classes and 20 super-classes of images. Instead of predicting the labels of fine-grained classes, we predict the labels of super-classes. We train a Wide-ResNet-28 on this classification task and get an accuracy of 73.33%, and utilize the pretrained model to initialize the AI classifier in both our proposed architecture and the baseline learning to defer method. We then compare our method with two alternative approaches:(a) Global, weighted output merging approach proposed in Section 4; (b) Consistent learning to defer estimator (Mozannar and Sontag 2020), where each time the decision is made by only one expert or AI model. We tune the hyper-parameter of learning rates for both settings with a set of $\{1e^{-3},3e^{-4},1e^{-4},3e^{-5},1e^{-5},3e^{-6},1e^{-6}\}$. The results of varying expert competence is given in Table 6.
@@
It is shown that with the same initialized AI classifier and the experts’ competences, our method outperforms the baseline L2D method significantly in the cases of effective human-AI collaboration. Global weighted merging method achieve very competitive results as $k$ gets larger. Similary as in Section 5.4.1, we further investigate the effective of varying non-expertise capability levels with expert competence level $k=5$. The results are shown in Figure 4 (b). We observe that our method has a greatest advantage over the baseline around $p=0.7$, and for other non-expertise capability levels, it is possible that the advantage is smaller than the advantage at $p=0.0$.
@@ 5.4.3 Results on HateSpeech
We also conducted experiments on Hate speech dataset, in which we apply the proposed architecture and baseline methods to determine if an instance is hatespeech, offensive language or neither. We follow (Mozannar and Sontag 2020) to create the synthetic experts for this task, which is: (a) If the tweet is in AAE then with probability $p$ the expert predict the correct label and otherwise predict uniformly at random. (b) If the tweet is not in AAE, the expert predict the correct label with probability $q$ and otherwise predict uniformly at random. Also, we consider three different expert probabilities $p$ and $q$, corresponding to a fair expert ($p=q=0.9$), a biased expert towards AAE ($p=0.75,q=0.9$), and a biased expert towards non AAE tweets ($p=0.9,q=0.75$). We use a pre-trained transformer encoder on Hatespeech to initialize our model as well as the classifier in the baseline model. We use a Adam optimizer with a learning rate of 1e-6 and a batch size of 32. The results are shown in Table 7. The collaborative method largely outperforms the model-only and human-only methods. Moreover, compared with the method of consistent estimator given in (Mozannar and Sontag 2020), our method achieves significant improvements for non-AAE biased and fair setting.
@@
5.4.4 Results on CIFAR10H and GalaxyZoo
Table 8: Results on CIFAR10H and GalaxyZoo with real human labels. * denote results of other methods reproduced by us.
CIFAR10H GalaxyZoo
Mean S.E. Mean S.E.
Confidence 95.09 0.40 81.83* 0.20
Consistent (Impute) 96.03 0.21 83.40* 0.19
Consistent (2-step) 96.29 0.25 81.97* 0.23
DiffTriage - - 83.80 0.35
Equal Weights 94.70 0.00 80.19 0.48
Pooled Weights 96.20 0.31 83.13 0.12
Flexible Weights 96.53 0.32 84.46 0.18
CIFAR10H and Galaxyzoo are image classification datasets with real human labels. For CIFAR-10H, we split it with 0.7:0.15:0.15 for training set, validation set and test set. For each example, we random sample an expert’s label in each epoch. For Galaxyzoo, we apply the same pre-processing method as in (Okati et al. 2021) with 0.6:0.2:0.2 train-val-test split ratio, and perform binary classification with the weighted combination of both human score vectors and model outputs. We compare the proposed methods with (Raghu et al. 2019), (Mozannar and Sontag 2020) and (Okati et al. 2021). A pretrained WideResNet-28 with the same setting as in (Mozannar and Sontag 2020) is applied for finetuning the AI model as well as the feature extractor for both CIFAR10H and GalaxyZoo datasets. For our method, we apply three settings: The first is using equal weights for both human and AI model (denoted by “Equal Wights”). The second is using trained global weights, which means we update the weights for every example in each iteration, and then take average of them to get shared global weights for human and AI. The global weights are applied for calculating the combination scores (denoted as “Pooled weights”). The third is flexible weights, which is using instance aware weights for calculating the combination scores (denoted as “flexible weights”). We apply an Adam optimizer with learning rate 1e-4 for the weights generation model, and 1e-5 for pretrained AI model and feature extractors. The effective batch size is set to be 128. The checkpoints with the minimum validation loss in 50 epochs of training are applied for testing.
The results are given in Table 8. We report both the average accuracy of 5 trials with different seeds and the standard error of the average. As not all related papers report the numerical results on both two dataset, we reproduce the rest by ourselves. We observe that for both two datasets, the flexible weights setting achieves the highest classification accuracy. Our method with flexible weights setting significantly outperforms the method in (Okati et al. 2021) on Galaxyzoo dataset, and the method in (Raghu et al. 2019) on CIFAR-10H. It achieves the higher average performance than (Mozannar and Sontag 2020) on CIFAR-10H although the difference is less significant. Our reproduced accuracy of DiffTriage on CIFAR10H is lower than model-only accuracy (<<93%), which may due to the lack of tuning or the ineffectiveness of this approach on the CIFAR-10H dataset (more categories than Galaxyzoo and hatespeech).
Figure 5: Diagram of accuracy under collaborative decision by AI models and multiple human experts.
@@ 5.4.4 Results on CIFAR10H and GalaxyZoo
To investigate the effectiveness of the proposed architecture for the scenario with multiple human experts, we assume that we have multiple human experts with different expertise sets with equal size, for which each expert can have only 1 or 2 expertise categories. We consider the collaborative decision making by one model and these human experts. Also, we consider the case of solely using these experts for collaborative decision making without collaboration with AI model. The experiments are done on CIFAR10. We adopt a pretrained ResNet-18 model as the classifier as well as the image representation extractor. A 6 layer transformer encoder with 8 heads and a embedding dimension of 512 is applied. An Adam optimizer with a learning rate of 1e-5, and a batch size of 64 are applied. The result of the case that each expert has only 1 expertise category is shown in Figure 5(a), where $s$ is the size of experts group. We observe that when the number of experts is small, the performance of both AI+experts system and experts-only system increase linearly. However, when the number of experts becomes large, the increasing trends diminishes slowly. This indicates that we need to apply heavy hyper-parameter tuning or use a more powerful weight generation model for handling the collaboration with a large number of experts. In the second experiment of multiple expert collaborative decision making. we assume that each expert has an expertise set with size 2, and the 5 experts have expertise sets $\{1,2\}$, $\{3,4\}$, $\{5,6\}$, $\{7,8\}$, $\{9,10\}$, respectively. The result of this case that is shown in Figure 5(b), where We observe that similar trend as in Figure 5(a). The propose method provide a significant improvement of the system accuracy as number of expert increases, but there is a diminishing marginal utility effect when the expert number is large.
@@ 7.1 Comparing with related works
The proposed architecture adopts a strategy of merging the output vectors from both AI models and human experts, rather than merely deciding whether to defer the task to a human. By applying a dynamical weighted combination of the output vectors from both human and AI agents, our method outperforms the major learning to defer methods. In contrast to the work by (Kerrigan et al. 2021), which also focuses on output vector combination, our method incorporates a representation learning mechanism instead of relying on a naive Bayesian estimator based on a confusion matrix. This allows our approach to seamlessly accommodate multiple experts in an end-to-end manner.
Our method also facilitates human-AI collaborative decision-making involving multiple AI and human agents. This is achieved through the learning of capability vectors in a uniform dimensional space. Related studies on human-AI collaborative decision-making with multiple experts typically employ different datasets or base models in their experiments. For instance, (Mao et al. 2024a) utilize ResNet-4 as the predictive AI model, while employing ResNet-10, ResNet-16, and ResNet-28 with increasing capacities as the human experts, focusing on the CIFAR-10 and SVHN image classification datasets. Additionally, Hemmer et al. (2022) investigate three distinct tasks, including Hate Speech detection, CIFAR-100 classification, and the NIH Dataset. As different papers apply different experimental settings, it is hard to make a complete comparison. Current experimental results convince us the advantage of our method can be extend to other settings.
@@ 8 Conclusion
In this study, we present a human-AI collaborative decision-making architecture based on capability vectors, applicable to both generative and discriminative decision tasks. By utilizing capability vectors, we can uniformly model the decision intelligence of human experts and AI models across the same set of dimensions. Additionally, we developed a learning-free global weighted merging method for collaborative decision-making involving one human expert and one AI model specifically for classification tasks, which can be applied as a baseline for check the effectiveness of instance-aware weighted merging methods. Our experiments, conducted on image classification and sentiment analysis tasks, demonstrate that the proposed method outperforms existing approaches across various human-AI collaborative scenarios. Experiments are performed using both simulated experts on CIFAR10/CIFAR100/Hatespeech and real experts on CIFAR-10H and GalaxyZoo, which demonstrate the advantage of proposed method over existing state-of-the-art models. We also show that for a widely used expert model in classification tasks, our method is more robust to the change of non-expertise capability level. In addition, the proposed method can be applied for multiple expert collaborative decision with or without AI decision makers.